OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6). y = x + 7 y = −3x + 3 y = x + 3 y = −3x + 7

OpenStudy (anonymous):

What do you know about the new line? Parallel lines have the same slope. You also have a given point. Slope -3 Point (−1, 6) Use Point-Slope form. \[y - y1 = m(x - x1) \]y - 6 = -3(x - (-1)) y - 6 = -3(x + 1) \[Distribute\]y - 6 = -3x - 3 \[Add\quad 6\]y = -3x + 3

OpenStudy (anonymous):

\[\Huge{\color{RED}{\mathbb{Welcome \quad To \quad OpenStudy!}}} \]

OpenStudy (anonymous):